OpenStudy (anonymous):

Arnold needs a 25% solution of nitric acid. He has 20 ml of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution? A. 20 ml B. 15 ml C. 25 ml D. 10 ml

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

what should i do 1st

jimthompson5910 (jim_thompson5910):

let x = amount of the 15% solution

jimthompson5910 (jim_thompson5910):

you should always start with what you know

jimthompson5910 (jim_thompson5910):

and declare what you don't and try to work towards getting it

jimthompson5910 (jim_thompson5910):

He has 20 ml of a 30% solution so he has 20*0.3 = 6 mL of pure acid

jimthompson5910 (jim_thompson5910):

he also has x mL of the 15% solution, so he also has an additional 0.15x mL of acid

jimthompson5910 (jim_thompson5910):

in total, he has 0.15x + 6 milliliters of pure acid

jimthompson5910 (jim_thompson5910):

he has 20 mL of the 30% solution he has x mL of the 15% solution so he has 20+x mL of solution total

jimthompson5910 (jim_thompson5910):

he wants 25% of this final solution (20+x mL) to be pure acid, so he wants 0.25(20+x) = 5 + 0.25x mL of pure acid

jimthompson5910 (jim_thompson5910):

if you go back, you'll see that he has 0.15x+6 mL of pure acid, so we can say 0.15x + 6 = 5 + 0.25x

jimthompson5910 (jim_thompson5910):

solve that for x to get your answer

OpenStudy (anonymous):

D!

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

once you get the setup down, it's not so bad